### v3

# Robust, fast, and efficient formation of stable tetratomic molecules from ultracold atoms via generalized stimulated Raman exact passage

Abstract
The study of the conversion of ultracold atoms into molecules has long remained a hot topic in atomic, molecular, and optical physics. However, most prior research has focused on diatomic molecules, with relatively scarce exploration of polyatomic molecules. Here we propose a two-step strategy for the formation of stable ultracold tetratomic molecules. We first suggest a generalized nonlinear stimulated Raman exact passage (STIREP) technique for the coherent conversion of ultracold atoms to tetratomic molecules, which is subsequently followed by a chainwise-STIREP technique to transfer the resulting molecules into a sufficiently stable ground state. Through systematic numerical analysis, we demonstrate that the proposed two-step strategy holds great potential for the robust, fast, and efficient formation of stable ultracold tetratomic molecules.
I Introduction
The successful creation of various ultracold diatomic molecules has stimulated growing interest in polyatomic molecules (UPMs) Rittenhouse and Sadeghpour (2010); Zeppenfeld et al. (2012); Kozyryev et al. (2017); Yang et al. (2022); Vilas et al. (2024); Chen et al. (2024). In contrast to diatomic molecules, UPMs offer additional complexity Carr et al. (2009); Küpper et al. (2009); Patterson et al. (2010), lending themselves to numerous interesting studies in, for instance, cold chemistry Karman et al. (2024), precision measurements Hutzler (2020) and the realization of exotic quantum phases Schmidt et al. (2022), to name a few. In order to tap their full potential, however, long-lived ensembles and full quantum state control serve as prerequisites Langen et al. (2024). Thus, the efficient conversion of ultracold atoms into stable UPMs constitutes an essential requirement, which hitherto remains challenging.
In the ongoing experiments, ultracold molecules are first prepared by association of pre-cooled atoms Chin et al. (2010), either via photo- or the magneto-association Dulieu and Gabbanini (2009); Chandler (2010); Quéméner and Julienne (2012); Zhao and Pan (2022). Subsequently, the formed molecules should be immediately transferred to absolute ground states, it is commonly achieved via the well-known Stimulated Raman adiabatic passage (STIRAP) Drummond et al. (2002); Lang et al. (2008) and its variants, such as multi-state STIRAP Danzl et al. (2010); Mark et al. (2009); Kuznetsova et al. (2008); Jing et al. (2008); Qian et al. (2010); Devolder et al. (2021). In principle, by manipulating the system to evolve along the adiabatic dark state, an efficient and robust state transfer can be obtained Bergmann et al. (1998); Král et al. (2007); Vitanov et al. (2017). Until now, STIRAP has been demonstrated successfully Bergmann et al. (2019), however, its application has been limited to the realm of diatomic molecules Winkler et al. (2007); Lang et al. (2008); Danzl et al. (2008); Aikawa et al. (2010); Ospelkaus et al. (2008); Feng et al. (2009); Liu et al. (2019a); Duda et al. (2023); Zamarski et al. (2025). Directly applying it to UPMs preparation remains challenging. To address this problem, researchers have developed STIRAP variants, for instance, Feshbach/Efimov resonance-assisted STIRAP, to support theoretical investigations into UPMs preparation Jing et al. (2007); Meng et al. (2009); Li and Peng (2011); Dou et al. (2012). Nevertheless, the successful application of such techniques is limited by the adiabatic criterion Zhou et al. (2010), which renders the systems vulnerable to environmental factors, e.g., collisions or decay to outside states.
In order to relax limitations such as adiabaticity related to the STIRAP-based scheme, a technique which is called shortcut-to-adiabaticity (STA) was proposed Chen et al. (2010). By leveraging the STA, the adiabatic process can be expedited while ensuring that the outcomes remain consistent. One particularly successful method for designing such fast routes is counterdiabatic driving Demirplak and Rice (2005, 2008); Berry (2009); Finžgar et al. (2025); Morawetz and Polkovnikov (2025); Dengis et al. (2025a, b), which allows one to construct a modification of an original Hamiltonian to eliminate unwanted nonadiabatic couplings Guéry-Odelin et al. (2019). Currently, STA is widely used in theory to realize fast and robust ultracold molecule formation Zhu and Chen (2021); Zhang and Dou (2021); Cao et al. (2023) and control the selective vibrational population transfer of molecules Masuda and Rice (2015a, b). In a recent study, Zhu etet alalZhu and Chen (2021) proposed a fast-forward scaling-assisted STIRAP method to study the conversion from atoms to diatomic molecules. In another study, Vitanov developed a shortcut multi-state STIRAP technique Vitanov (2020), which may provide an alternative ways for forming stable ground-state ultracold molecules. Compared to the traditional three-level system, the employment of multi-level ones exhibits unique advantages; to illustrate, this strategy proves instrumental in resolving the challenge of weak Franck-Condon factors (FCFs) during the process of both atom-to-molecule conversion Jaksch et al. (2002); Qian et al. (2010) and ground-state molecular transfer Danzl et al. (2010); Kuznetsova et al. (2008). Despite these advances, the major problem is that most prior studies usually require introducing additional couplings for eliminating non-adiabatic effects, which may pose additional experimental challenges. On the other hand, in 2017, Dorier etet alalDorier et al. (2017) suggested an alternative method, namely nonlinear stimulated Raman exact passage (STIREP), for the efficient atom-to-molecule conversion. This method conforms to the spirit of STA, because the control fields are determined by the specific dynamics selected to achieve exact state transfer Laforgue et al. (2022); Dridi et al. (2024). This trick has been extended to the formation of triatomic molecules Cao (2024). Currently, an important question in this realm is whether it is possible to extend this successful tool to form larger polyatomic molecular complexes Dou et al. (2013); Shammout et al. (2025); Quéméner et al. (2023); Cao et al. (2024); Gacesa et al. (2021); Gu et al. (2025), and generalize it to transfer of molecules into deeply-bound states Koch and Moszyński (2008); Danzl et al. (2010); Kuznetsova et al. (2008); de Lima (2024); Mukherjee and Tomza (2025).
The purpose of this paper is to develop a theoretical formalism that extends the existing STIREP protocols to the case that permits control of conversion from atoms to stable tetratomic molecules. We first report a generalization of the STIREP technique for controlling the conversion from ultracold atoms to tetratomic molecules. To transfer the resulting molecules to a sufficiently stable ground state in a short time, we further develop a chainwise-STIREP (abbreviated as C-STIREP hereafter) technique. We present the complete explanations of the underlying mechanism and the corresponding numerical analyses. The results demonstrate that our proposed two-step strategy performs well in both the preparation of ultracold tetratomic molecules and the ground-state molecular transfer.
This paper is outlined as follows. In Sec. II, we formulate the model for the conversion of atoms to tetratomic molecules, followed by a discussion on the underlying mechanism of the generalized N-STIREP technique. Numerical analyses are presented. In Sec. III, we suggest and analysis a chainwise-STIREP for molecular transfer into stable ground state based on a five-state molecular model. Section. IV concludes the paper and presents an outlook.
II COHERENT CONVERSION FROM ATOMS TO TETRATOMIC MOLECULES VIA GENERALIZED NONLINEAR STIREP
II.1 Model and Method
As outlined in the introduction, our top priority is to prepare homonuclear tetratomic molecules. Figure. 1(a) shows a visualization of the process.
Figure 1: (Color online) Schematic showing the coherent conversion from ultracold atoms to tetratomic molecules.
Starting from a Bose condensate of atoms, we apply the field $\Omega_{1}^{\prime}$ to associate free atoms $|a\rangle$ into triatomic molecules $|m\rangle$; the other field $\Omega_{2}^{\prime}$, in turn, is responsible for associating pairs composed of an atom and a triatomic molecule $|m\rangle$ into tetratomic molecules $|t\rangle$. For the applied lasers, their frequencies are expressed using the single-photon and two-photon detunings $\delta$ and $\Delta$, respectively. We remark that the formation of triatomic molecules via the photoassociation of free atoms has been recently reported Schnabel et al. (2021), whereas the formation of tetratomic molecules through the coupling of an atom-triatomic molecule pair has also been investigated Jing and Jiang (2008). Our aim is to find an exact solution for controlling fast and efficient conversion from ultracold atoms to tetratomic molecules by generalizing existing N-STIREP protocols.
The Hamiltonian of this system, in the interaction picture with $\hbar=1$, takes the form of
$\begin{aligned} \hat{\mathcal{H}}\!=&-\int\!d\mathbf{r}\left\{\sum_{ij}\chi_{ij}^{\prime}\hat{\psi}_{i}^{\dagger}(\mathbf{r})\hat{\psi}_{j}^{\dagger}(\mathbf{r})\hat{\psi}_{j}(\mathbf{r})\hat{\psi}_{i}(\mathbf{r})\!+\!\delta\hat{\psi}_{\mathrm{m}}^{\dagger}(\mathbf{r})\hat{\psi}_{\mathrm{m}}(\mathbf{r})\right.\\ &\!+\!\Omega_{1}^{\prime}\{\hat{\psi}_{\mathrm{m}}^{\dagger}(\mathbf{r})[\hat{\psi}_{\mathrm{a}}(\mathbf{r})]^{3}+\mathrm{H.c.}\}\!+\!(\Delta\!+\!\delta)\hat{\psi}_{\mathrm{t}}^{\dagger}(\mathbf{r})\hat{\psi}_{\mathrm{t}}(\mathbf{r})\\ &\!-\!\Omega_{2}^{\prime}[\hat{\psi}_{\mathrm{t}}^{\dagger}(\mathbf{r})\hat{\psi}_{\mathrm{m}}(\mathbf{r})\hat{\psi}_{\mathrm{a}}(\mathbf{r})+\mathrm{H.c.}]\Bigg\}.\end{aligned}$ (1)
where $\hat{\psi}_{k}^{\dagger}(\mathbf{r})$ and $\hat{\psi}_{k}^{\dagger}(\mathbf{r})$ are, respectively, the bosonic creation and annihilation operators for state $|k\rangle(k\!=\!a,m,t)$. The indices $i,j=a,m,t$ represent the atomic, triatomic molecular, and tetratomic molecular states, respectively. By adopting the mean-field approach $\hat{\psi}_{i}\rightarrow\sqrt{n}{\psi}_{i}$ ($n$ denotes the initial atomic density) Ling et al. (2004), the Heisenberg motional equation can be described as
$\begin{aligned} i\dot{\psi}_{a}\!&=\!-2K_{a}\psi_{a}\!-\!3\Omega_{1}{\psi_{m}\psi^{\ast}_{a}}^{2}\!+\!\Omega_{2}\psi_{t}{\psi^{*}_{m}},\\ i\dot{\psi}_{m}\!&=\!-2K_{m}\psi_{m}\!-\!(\delta\!+\!i\gamma_{m})\psi_{m}\!-\!\Omega_{1}\psi_{a}^{3}\!+\!\Omega_{2}\psi_{t}{\psi^{\ast}_{a}},\\ i\dot{\psi}_{t}\!&=\!-2K_{t}\psi_{t}\!-\!(\Delta\!+\!\delta\!+\!i\gamma_{t})\psi_{t}\!+\!\Omega_{2}\psi_{m}\psi_{a}.\end{aligned}$ (2)
In the above equations, $\chi_{ij}\!=\!n\chi_{ij}^{\prime},\Omega_{1}\!=\!n\Omega_{1}^{\prime},$ and $\Omega_{2}\!=\!\sqrt{n}\Omega_{2}^{\prime}$ denote the renormalized quantities. $\gamma_{m}$ and $\gamma_{t}$ are introduced to phenomenologically describe the losses of the corresponding states to other undetected states. Nevertheless, we temporarily ignore these losses in the theoretical analysis to clarify the underlying physics. In this case, we have $|\psi_{\mathrm{a}}|^{2}\!+\!3|\psi_{\mathrm{m}}|^{2}\!+\!4|\psi_{\mathrm{t}}|^{2}\!=\!1$. The third-order nonlinearity term $K_{i}\!=\!\sum_{j}\chi_{ij}\left|\psi_{j}\right|^{2}$ characterizes elastic collisions between particles, which gives rise to the dynamical instability effect Itin and Watanabe (2007); Guérin et al. (2013).
In principle, the effects of third-order nonlinearity can be eliminated by employing resonance locked inverse engineering Dorier et al. (2017); Zhu et al. (2023); Cao et al. (2023). To this end, a general parametrization with the phase $\alpha$ should be introduced, i.e., $\psi_{\mathrm{a}}\!\mapsto\!\psi_{\mathrm{a}}e^{i\alpha},\psi_{\mathrm{m}}\!\mapsto\!\psi_{\mathrm{m}}e^{\mathrm{3}i\alpha},\psi_{\mathrm{t}}\!\mapsto\!\psi_{\mathrm{t}}e^{\mathrm{4}i\alpha}$. From this, the condition for the resonance locking can be derived as
$\begin{aligned} \delta=6K_{\mathrm{a}}\!-\!2K_{\mathrm{m}},\quad\Delta=2(K_{\mathrm{a}}\!+\!K_{\mathrm{m}}\!-\!K_{\mathrm{t}}).\end{aligned}$ (3)
With the help of condition (3), the system of equations (2) becomes
$\begin{aligned} i\dot{\psi}_{\mathrm{a}}\!&=\!-\Omega_{1}{\psi_{\mathrm{m}}\psi^{\ast}_{\mathrm{a}}}^{2}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{*}_{\mathrm{m}}},\\ i\dot{\psi}_{\mathrm{m}}\!&=\!-\Omega_{1}\psi_{\mathrm{a}}^{3}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{\ast}_{\mathrm{a}}},\\ i\dot{\psi}_{\mathrm{t}}\!&=\!\Omega_{2}\psi_{\mathrm{m}}\psi_{\mathrm{a}}.\end{aligned}$ (4)
Now two optical fields $\Omega_{1,2}$ are resonant with the transitions $|\mathrm{a}\rangle\leftrightarrow|\mathrm{m}\rangle$ and $|\mathrm{m}\rangle\leftrightarrow|\mathrm{t}\rangle$, respectively.
We further note that, in our model, the couplings of both fields are nonlinear, which differs from Ref. Dorier et al. (2017), where only the coupling of the first field is nonlinear. In what follows, we demonstrate how the generalized STIREP adapts to the system to compensate for the nonlinearities induced by these couplings. To this end, we introduce two dynamical angles $\theta$ and $\phi$ to parameterize the state vector:
$\begin{aligned} &\psi_{\mathrm{a}}\!=\!\cos\phi\cos\theta,\\ &\psi_{\mathrm{m}}\!=\!-i\frac{\sin\phi}{\sqrt{3}},\\ &\psi_{\mathrm{t}}\!=\!-\frac{\cos\phi\sin\theta}{2}.\end{aligned}$ (5)
Inserting this into Eqs. (4) we can get
$\begin{aligned} \Omega_{1}&\!=\!-\frac{\sqrt{3}}{12}\left(\frac{1\!+\!3\cos^{2}\theta}{\cos^{3}\theta\cos^{2}\phi}\dot{\phi}\!+\!\frac{10\tan\theta\dot{\theta}}{\sin 2\phi\cos\theta}\right),\\ \Omega_{2}&\!=\!\frac{\sqrt{3}}{2}\left(\frac{\dot{\theta}}{\sin\phi}\!-\!\frac{\dot{\phi}\tan\theta}{\cos\phi}\right).\end{aligned}$ (6)
To ensure the desired atom-to-tetratomic molecular conversion, we apply boundary conditions: $0\!\leftarrow\!\phi\!\rightarrow\!0$ and $0\!\leftarrow\!\theta\!\rightarrow\!\pi/2$, where the arrows to the left and right signify the limits when $t\!\rightarrow\!t_{i}$ and $t\!\rightarrow\!t_{f}$, respectively. The controlled parameters $\theta$ and $\phi$ are chosen as
$\begin{aligned} \theta(t)&=\frac{\pi}{4}\eta\left(\mathrm{tanh}\left(\frac{t}{T}\right)\!+\!1\right),\\ \phi(t)&=\frac{4\epsilon\sqrt{\theta(t)\left[\frac{\pi}{2}\!-\!\theta(t)\right]}}{\pi},\end{aligned}$ (7)
where $\epsilon$ and $\eta$ are constants. $\epsilon\gtrsim 0$ enables the regulation of the transient population in the intermediate state, while $\eta\gtrsim 1$ governs the final conversion efficiency. It should be noted that $\eta=1$ is prohibited in nonlinear cases, whereas no such restriction exists in linear cases.
II.2 Result and Discussion
Figure 2: (Color online) Rabi frequencies (left column) and populations (left column) as a function of $t$. Adopted Parameters: $\eta=0.96,\epsilon=0.2$.
To demonstrate the validity of our procedure, we are going to carry out numerical calculations on population dynamics by employing the standard fourth-order Runge-Kutta method.
Figure. 2(a) shows the time-dependent evolution of Rabi frequencies used for the conversion of atoms to tetratomic molecules. In Fig. 2(b), we present the corresponding population dynamics. As anticipated, we achieved an almost perfect conversion efficiency, while maintaining a small transient population in the intermediate state during the interaction. Clearly, our method has the potential to maintain the phase-space density of ultracold mixtures, as a high conversion efficiency is essential for this maintenance.
To better illustrate the advantages of our protocol, we simulated its performance and compared it with that of the conventional STIRAP protocol. This study is feasible, because the transfer process we consider can be viewed as an abstract three-level model. In our simulations, the STIRAP process was implemented using a pair of Gaussian pulses Meng et al. (2008),
[…]
where $T$, $t_{\mathrm{P,S}}$ and $\Omega^{0}_{\mathrm{P,S}}$ denote the width, peak times and peak strengths of the Gaussian pulses, respectively. The Stokes pulse $\Omega_{\mathrm{S}}$ is applied first to induce the transition $|m\rangle\leftrightarrow|t\rangle$ and is followed by the pump pulse $\Omega_{\mathrm{P}}$, which induces the transition $|a\rangle\leftrightarrow|m\rangle$.
Figure 3: (Color online) (a) Rabi frequencies as functions of $t$ under the generalized STIREP and the STIRAP. $\chi=8.4437$; (b) Conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$ under the (upper surface) generalized STIREP and (lower surface) STIRAP protocols; (c) Conversion efficiency as a function of $T$ under the generalized STIREP; Parameters: $\gamma_{m}=10^{-1}/T$, $\gamma_{t}=10^{-4}/T$.
Figure 4: (Color online) (a) Contour plot showing conversion efficiency versus $\delta T$ and $\delta\Omega_{1}$, where $\Omega_{2}$ is fixed at its ideal value without any deviations; (b) Contour plot showing conversion efficiency versus $\delta T$ and $\delta\Omega_{2}$, where $\Omega_{1}$ is fixed at its ideal value. Note that parameter deviations are implemented on the basis of the original parameters in Fig. 2.
To ensure a fair comparison, both schemes are carried out with the same evolution time, peak times, and peak strengths. According to Fig. 3(a), we set $t_{\mathrm{P}}\!=\!1.727T$, $t_{\mathrm{S}}\!=\!0.006T$, $\Omega^{0}_{\mathrm{P}}=27.7204/T$, and $\Omega^{0}_{\mathrm{S}}=3.283/T$, respectively. Meanwhile, the robustness of both schemes against Rabi frequency errors will be assessed. To this end, we systematically modify the amplitudes of the original Rabi frequencies while preserving their waveform shapes, with specific adjustments defined as $\Omega_{\mathrm{1,P}}\rightarrow(1-\eta_{1})\Omega_{\mathrm{1,P}}$ and $\Omega_{\mathrm{2,S}}\rightarrow(1-\eta_{2})\Omega_{\mathrm{2,S}}$, respectively.
Figure. 3(a) displays the time sequence of the Rabi frequencies corresponding to our protocol and STIRAP. It is obvious that, like the Gaussian waveform, the fields used in our protocol are still sufficiently smooth. We plot in Fig. 3(b) the conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$. Here, the upper surface corresponds to our protocol, whereas the lower surface corresponds to STIRAP. As it is observed, our protocol always performs better than STIRAP throughout the whole range of error parameters. In addition, we stress that our protocol allows for faster control over the conversion process. We demonstrate this fact by plotting a curve of the conversion efficiency of our protocol as a function of the evolution time $T$, as depicted in Fig. 3(c). As it turns out, the higher conversion efficiency can be achieved by shortening the evolution time. This intriguing phenomenon aligns with the core spirit of STA.
Further studies can reveal more properties of our protocol. In Fig. 4(a), by fixing $\Omega_{2}$ at its ideal value, we present a contour plot of transfer efficiency against the evolution time deviation $\delta T$ and the Rabi frequency deviation $\delta\Omega_{1}$. As a comparison, we study in Fig. 4(b) transfer efficiency against $\delta T$ and $\delta\Omega_{2}$, with $\Omega_{1}$ fixed at its ideal value. It should be noted that the values marked by the white dashed lines in these figures are equal to the ideal value under the working condition of no deviations; this value correspond exactly to the end value of the red curve in Fig. 2(b), and its specific value is $0.9899$. It is observed that even for deviations of $|\delta T/T|\!=\!|\delta\Omega_{1}/\Omega_{1}|\!=\!|\delta\Omega_{2}/\Omega_{2}|\!=\!10\%$, the conversion efficiencies remain remarkably close to the ideal value. The results further demonstrate that the C-STIREP exhibits high robustness with respect to imperfections in operation.
The above calculations demonstrate that the robustness and fast conversion of our protocol facilitate the preparation of ultracold tetratomic molecules. Nevertheless, it should be stressed that the resulting molecules are usually energetically unstable Kuznetsova et al. (2009), thereby making it highly necessary from a practical standpoint to rapidly transfer them to deeply-bound vibrational states. To this end, we suggest a five-state C-STIREP technique, whose details will be further analyzed and presented in the following section.
III Ground-state molecular population transfer via C-STIREP
III.1 Model and Method
In this section, our goal is to generalize the STIREP into multi-state system, with the motivation to transfer the weakly-bound molecules into a sufficiently stable state. Inspired by the double-STIRAP transfer scheme Danzl et al. (2010); Kuznetsova et al. (2008), we restrict ourselves to a five-level chainwise-coupled molecular condensate system. As presented in Fig. 5, all five molecular states are coupled in a distorted M-type configuration,
Figure 5: (Color online) A schematic of coherent population transfer from weakly- to deeply-bound molecules.
in which four fields with Rabi frequencies $\Omega_{i}(i=1,2,3,4)$ are applied to move the molecules from a high vibrational level of the ground electronic state $|c\rangle$ (now relabeled as $|1\rangle$) to a sufficiently stable vibrational level of the ground electronic state $|5\rangle$. In the entire transfer process, three intermediate states are introduced to act as the “bridge”, where $|3\rangle$ is regarded as an intermediate vibrational level of ground electronic state, $|2\rangle$ and $|4\rangle$ represent two vibrational levels of the excited electronic state. To render this problem tractable, we shall analyze this problem under the collisionless limit, as done in prior Refs. Pu et al. (2007); Jing et al. (2008); Qian et al. (2010); Graefe et al. (2015). Accordingly, the Hamiltonian of the system can be described in the following form:
$\begin{aligned} \hat{\mathcal{H}}\!=\!\sum_{i=2}^{5}\Delta_{i}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i}\!+\!\sum_{i=1}^{4}(\Omega_{i}\hat{\psi}_{i}^{\dagger}\hat{\psi}_{i+1}\!+\!\mathrm{H.c.}).\end{aligned}$ (9)
The Heisenberg motional equation can be written as
$\begin{aligned} i\dot{\psi}_{1}&\!=\!\Omega_{1}\psi_{2},\\ i\dot{\psi}_{2}&\!=\!\Omega_{1}\psi_{1}\!+\!\Delta_{1}\psi_{2}\!+\!\Omega_{2}\psi_{3},\\ i\dot{\psi}_{3}&\!=\!\Omega_{2}\psi_{2}\!+\!\Omega_{3}\psi_{4},\\ i\dot{\psi}_{4}&\!=\!\Omega_{3}\psi_{3}\!+\!\Delta_{2}\psi_{4}\!+\!\Omega_{4}\psi_{5},\\ i\dot{\psi}_{5}&\!=\!\Omega_{4}\psi_{4}.\end{aligned}$ (10)
For simplicity, we have assumed that $\Delta_{3}\!=\Delta_{5}\!=\!0$, and $\Delta_{2}\!=\!\Delta_{4}\!=\!\Delta$. In what follows, we elaborate on how the C-STIREP technique operates within the above system.
First, we postulate that $\Delta$ is the largest evolution frequency Lugiato et al. (1984); Mackie and Phou (2010); Liu et al. (2019b), $\Delta\!\gg\!\Omega_{i}(i\!=\!1,2,3,4)$. In this limit, adiabatic elimination (AE) enables the derivation of an effective three-state model, which takes the form of
$\begin{aligned} i\dot{\psi}_{1}&=\varepsilon_{11}\psi_{1}\!+\!\omega_{13}\psi_{3},\\ i\dot{\psi}_{3}&=\varepsilon_{22}\psi_{3}\!+\!\omega_{13}\psi_{1}\!+\!\omega_{35}\psi_{5},\\ i\dot{\psi}_{5}&=\varepsilon_{33}\psi_{5}\!+\!\omega_{35}\psi_{3}.\end{aligned}$ (11)
In which, the effective Rabi frequencies are $\omega_{13}=\!-\!\frac{\Omega_{1}\Omega_{2}}{\Delta}$ and $\omega_{35}=\!-\!\frac{\Omega_{3}\Omega_{4}}{\Delta}$; the effective detunings are $\varepsilon_{11}=\!-\!\frac{\Omega_{1}^{2}}{\Delta},\varepsilon_{22}=\!-\!\frac{\Omega_{2}^{2}+\Omega_{3}^{2}}{\Delta}$ and $\varepsilon_{33}=\!-\!\frac{\Omega_{4}^{2}}{\Delta}$, respectively. To continue, let us set $\varepsilon_{11}=\varepsilon_{22}=\varepsilon_{33}=\varepsilon_{0}$, from which the condition for the generalized resonance locking follows,
$\Omega_{1}(t)=\Omega_{4}(t)=\sqrt{\Omega_{2}^{2}(t)+\Omega_{3}^{2}(t)}.$ (12)
By further performing the simple transformation $\psi_{j}\!=\!\psi^{{}^{\prime}}_{j}e^{-i\varepsilon_{0}t}(j\!=\!1,3,5\!)$, we can obtain
$\begin{aligned} i\dot{\psi}^{{}^{\prime}}_{1}&\!=\!\omega_{13}\psi^{{}^{\prime}}_{3},\\ i\dot{\psi}^{{}^{\prime}}_{3}&\!=\!\omega_{13}\psi^{{}^{\prime}}_{1}\!+\!\omega_{35}\psi^{{}^{\prime}}_{5},\\ i\dot{\psi}^{{}^{\prime}}_{5}&\!=\!\omega_{35}\psi^{{}^{\prime}}_{3},\end{aligned}$ (13)
in which
$\begin{aligned} \omega_{13}\!=-\!\frac{\Omega_{2}\sqrt{\Omega_{2}^{2}\!+\!\Omega_{3}^{2}}}{\Delta},\quad\omega_{35}\!=-\!\frac{\Omega_{3}\sqrt{\Omega_{3}^{2}\!+\!\Omega_{2}^{2}}}{\Delta}.\end{aligned}$ (14)
Having obtaining the equivalent model (13), we immediately parameterize the state vector as
$\begin{aligned} &\psi^{{}^{\prime}}_{1}\!=\!\cos\phi\cos\theta,\\ &\psi^{{}^{\prime}}_{3}\!=\!-i\sin\phi,\\ &\psi^{{}^{\prime}}_{5}\!=\!-\cos\phi\sin\theta,\end{aligned}$ (15)
The Rabi frequencies can be constructed as
$\begin{aligned} \omega_{13}&=\dot{\theta}\cot\phi\sin\theta\!+\!\dot{\phi}\cos\theta,\\ \omega_{35}&=\dot{\theta}\cot\phi\cos\theta\!-\!\dot{\phi}\sin\theta.\end{aligned}$ (16)
In order to drive the state transfer of interest, we apply boundary conditions $0\!\leftarrow\!\phi\!\rightarrow\!0,0\!\leftarrow\!\theta\!\rightarrow\!\pi/2$.
Up to now, we have obtained the STIREP for the equivalent system (13). In principle, by means of the two fields given by Eq. (16), one could leverage the advantages of STIREP for achieving the desired population transfer; however, it should be noted that such two-photon transitions are likely to be challenging to induce directly in practice. Thus, the remaining task is to engineer physically feasible driving fields within the five-level framework.
Now we are ready to design the modified Rabi frequencies for the original system. Like Eq. (14), we impose the following form on the two middle Rabi frequencies:
$\begin{aligned} \omega_{13}=-\frac{\tilde{\Omega}_{2}\sqrt{\tilde{\Omega}_{2}^{2}\!+\!\tilde{\Omega}_{3}^{2}}}{\Delta},\quad\omega_{35}=-\frac{\tilde{\Omega}_{3}\sqrt{\tilde{\Omega}_{2}^{2}\!+\!\tilde{\Omega}_{3}^{2}}}{\Delta}.\end{aligned}$ (17)
Meanwhile, we invoke the resonance locking condition (12) to modify the remaining two Rabi frequencies:
$\begin{aligned} \tilde{\Omega}_{1}=\tilde{\Omega}_{4}=\sqrt{\tilde{\Omega}_{2}^{2}+\tilde{\Omega}_{3}^{2}}.\end{aligned}$ (18)
Solving Eqs. (17) and (18), we thus obtain
$\begin{aligned} \tilde{\Omega}_{2}&\!=\!\left(\frac{\Delta^{2}\omega^{4}_{13}}{\omega^{2}_{13}\!+\!\omega^{2}_{35}}\right)^{\frac{1}{4}},\quad\tilde{\Omega}_{3}\!=\!\left(\frac{\Delta^{2}\omega^{4}_{35}}{\omega^{2}_{13}\!+\!\omega^{2}_{35}}\right)^{\frac{1}{4}},\\ \tilde{\Omega}_{1}&\!=\!\tilde{\Omega}_{4}\!=\!\left[\Delta^{2}(\omega^{2}_{13}\!+\!\omega^{2}_{35})\right]^{\frac{1}{4}}.\end{aligned}$ (19)
With the newly designed Rabi frequencies, we thereby develop a five-state C-STIREP technique. It should be noted that, to ensure the validity of this method, the AE condition $\Delta\!\gg\!\tilde{\Omega}_{i}(i\!=\!1,2,3,4)$ should still be satisfied Li and Chen (2016).
III.2 Result and Discussion
Now we numerically investigate the coherent population dynamics of the C-STIREP method. Unlike the above (7), we demonstrate that the enhanced performance of C-STIREP can also be achieved by using other pulse shapes. As an example, we consider
$\begin{aligned} \phi(t)&=\frac{\beta}{2}\left[1-\cos(\frac{2\pi t}{T})\right],\\ \theta(t)&=\frac{\pi t}{2T}-\frac{1}{3}\sin(\frac{2\pi t}{T})+\frac{1}{24}\sin(\frac{4\pi t}{T}).\end{aligned}$ (20)
Once $\phi$ and $\theta$ are determined, the effective Rabi frequencies can be found according to Eq. (16).
Figure 6: (Color online) (a) The sequence of four Rabi frequencies and (b) the population evolution of the five-state molecular system under the C-STIRAP. Adopted parameters: $\Delta\!=\!2500/T,\beta\!=\!0.25\pi$.
Figure. 6(a) illustrates the Rabi frequencies required to achieve the desired population transfer result. From this figure, it can be observed that the first and fourth Rabi frequencies not only exhibit the same temporal profile and equal magnitudes, but also straddle the middle two ones. They feature near-Gaussian profiles with sufficient smoothness, thus posing no actual implementation challenges Du et al. (2016); Zheng et al. (2022); Norambuena et al. (2024). The time evolution results from the initial state $|1\rangle$ to the target state $|5\rangle$ is displayed in Fig. 6(b). It can be observed that a complete state transfer can be achieved by employing four modified fields. A high transfer efficiency is helpful to minimize the loss in phase-space density when an ensemble of weakly-bound molecules is transferred to the ground state. In addition, we note that, throughout the entire transfer process, the intermediate state $|3\rangle$ receives some amount of transient population, while the populations of the two excited states are negligible. This process does not rely on the existence of dark states; instead, we utilize the AE protocol to achieve the stable decoupling of two excited states from dynamics. Thus, although these two states are prone to decaying out of the system, the transfer process remains scarcely swayed by them.
Figure 7: (Color online) (a) The transient population of state $|3\rangle$ as a function of $\beta$, where the other parameters are consistent with those in Fig. 6; (b) The populations of state $|3\rangle$ and state $|5\rangle$ as a function of $T$, where $\beta$ is fixed at $0.1\pi$, and all other parameters are consistent with those in Fig. 6.
Once a desired population transfer from $|1\rangle$ to $|5\rangle$ is achieved, we would like to impose further conditions of minimizing the amount of transient population that is in $|3\rangle$ throughout the process. Since the dynamics of the five-state system is equivalent to that of the reduced three-state one, the transient population of the intermediate state $|3\rangle$ can be accurately derived as $\sin^{2}\phi$. On this basis, we may attempt to reduce the value of $\phi$ to diminish the population of this state. We plot in Fig. 7(a) the transient population of state $|3\rangle$ as a function of $\beta$; please note that all the other parameters are maintained the same as those in Fig. 6. Clearly, the population of the intermediate state $|3\rangle$ witnesses a substantial decrease as the value of $\beta$ is gradually reduced. However, the population of this state never equals 0 since the limit of $\beta\rightarrow 0$ cannot be achieved in practical scenarios [see Eqs. (19) and (20)]. From the preceding analysis, it can be concluded that the five-state transfer scheme offers an additional benefit over the three-state counterpart: the former only requires us to control the transient population of the intermediate state $|3\rangle$, while the latter necessitates the control of the short-lived excited state. Given that the intermediate state $|3\rangle$ typically exhibits a sufficiently long lifetime, even a small amount of transient population within this state exerts a negligible influence on the final transfer efficiency; in contrast, the transfer efficiency of the three-state scheme is largely constrained by the characteristics of its intermediate short-lived excited state Ciamei et al. (2017); Li et al. (2017).
Figure 8: (Color online) (a) Contour plot showing transfer efficiency versus $\delta\tilde{\Omega}_{1}$ and $\delta\tilde{\Omega}_{4}$, where $\Omega_{2}$ and $\Omega_{3}$ are set to their ideal values without any deviation; (b) Contour plot showing transfer efficiency versus $\delta\tilde{\Omega}_{2}$ and $\delta\tilde{\Omega}_{3}$, where $\Omega_{2}$ and $\Omega_{3}$ passively vary according to condition (18). The parameter deviations are implemented on the basis of the adopted parameters in Fig. 7(b).
Furthermore, we point out that the transient population of the intermediate state $|3\rangle$ can be further reduced by leveraging the speed advantage of C-STIREP. To verify this finding, we plot in Fig. 7(b) the variation of populations of state $|3\rangle$ and state $|5\rangle$ with the evolution time $T$. As can be seen from this figure, the proposed method is capable of significantly shortening the population transfer time without deteriorating the transfer efficiency. This interesting feature mitigates the interaction duration between the system and its surrounding environment, thereby enhancing the transfer efficiency for practical applications.
Finally, we are going to perform the targeted numerical simulations to verify the necessity of Condition (18). Figure. 8(a) shows a contour plot of transfer efficiency against the Rabi frequency deviations $\delta\tilde{\Omega}_{1}$ and $\delta\tilde{\Omega}_{4}$ when $\tilde{\Omega}_{2}$ and $\tilde{\Omega}_{3}$ are fixed at their ideal values. Clearly, $\delta\!=\!0$ implies that an ideal operating condition is achieved; $\delta\!\neq\!0$ indicates that the Condition (18) is violated. According to the results shown in this figure, we can find that any deviation from the ideal condition will result in a significant drop in efficiency. In particular, when the deviation is relatively large, the transfer efficiency will suffer a loss of around $10\%$. Except for this, we can observe that the deviations $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ exert symmetric effects on the transfer efficiency. This confirms that this condition is an indispensable prerequisite for obtaining high population transfer efficiency. The transfer efficiency can remain at a high level if and only if $\tilde{\Omega}_{1}\!=\!\tilde{\Omega}_{4}$ and they should exactly match the ideal value.
On the other hand, Figure. 8(b) presents a contour plot of the transfer efficiency as a function of $\delta\tilde{\Omega}_{2}$ and $\delta\tilde{\Omega}_{3}$, with Condition (18) maintained; in this case, $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ will passively vary with the other two ones. It is clear that the transfer efficiency can keep at a high level, even when the deviation of the Rabi frequency deviations within $|\delta\tilde{\Omega}_{2}/\tilde{\Omega}_{2}|\!=\!|\delta\tilde{\Omega}_{3}/\tilde{\Omega}_{3}|\!=\!10\%$. This result further demonstrates the necessity of Condition (18).
IV Conclusions and Outlooks
To conclude, we present a theoretical formalism for the formation of stable ultracold tetratomic molecules. The process begins with a mixture of ultracold atoms: first, generalized nonlinear STIREP is employed to coherently assemble atoms into tetratomic molecules; subsequently, to ensure the resulting molecules exhibit sufficient stability, we propose that the C-STIREP technique be used to transfer them to the stable ground state. Numerical analyses show that the proposed two-step strategy enables the production of stable ultracold molecules with nearly ideal efficiency, while featuring excellent accelerability and high robustness.
The main advantages of our work are organized as follows: (i) It paves a potential pathway for forming stable ultracold tetratomic molecules, and in principle, has the potential to be extended to preparing larger polyatomic molecular complexes Dou et al. (2013); (ii) Unlike traditional methods, the adiabatic condition is released in our approach, thus enabling fast speed and high efficiency. Additionally, the transient populations of short-lived states can be effectively suppressed without relying on the dark state; (iii) Although the protocol follows the spirit of STA, it does not introduce extra couplings, and the applied pulses pose no actual implementation difficulties; (iv) Numerous exact solutions support the achievement of predefined goals, and the above-discussed examples (7) and (20) are not exhaustive Chen and Muga (2012); Kang et al. (2016); Baksic et al. (2016); Wu et al. (2019).
Finally, we point out that the tetratomic molecule preparation process suggested in this study aligns closely with the standard procedure for diatomic molecule, with both following the similar route of “first creating unstable molecular samples by assembling pre-cooled atom, and then converting them to sufficiently stable ground state”. Inspired by the studies in Refs. Mackie and Phou (2010); Price and Yelin (2019), we will, in the future, commit to exploring a coherent strategy that starts directly from ultracold atoms and enables the robust, fast and efficient production of stable UPMs. This problem may be worth mentioning.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants No. 12475026, No. 12075193 and No. 12365004) and the Doctoral Research Startup Fund of Northwest Normal University (Grant No. 6014/202503101301).